OpenStudy (anonymous):

turning (quadratic form) " 2x^2+8x+7=0 " into vertex form

OpenStudy (ranga):

2x^2+8x+7 = 2(x^2 + 4x + 7/2) Complete the square for the expression inside the parenthesis.

OpenStudy (ranga):

2x^2+8x+7 = 2(x^2 + 4x + 7/2) = 2{ (x+2)^2 - 4 + 7/2 } = 2{ (x+2)^2 - 1/2 } = 2(x+2)^2 - 1 2(x+2)^2 - 1 is the vertex form. The vertex of the parabola is is at (-2,-1)

OpenStudy (anonymous):

:D that was my answer i just solved it! thank you so much!!

OpenStudy (ranga):

You are welcome.